# Pseudo-Legendrian and Legendrian Simplicity of Links in 3-Manifolds

### v1

@@ 5. Examples
Let $V_{k}$ be a co-orienting vector fied of $\xi_{k}$. Let $K_{1}\sqcup K_{2}$ be a $V_{k}$-transverse link in $M$. Let $L=L_{1}\sqcup L_{2}$ be a Legendrian representative of the smooth isotopy class of $K_{1}\sqcup K_{2}$, obtained by taking a $C^{0}$-approximation, is in [etnyre2005legendrian]. Consider integers $i_{1},i_{2}\in\mathbb{Z}$. Recall from Section 2 that, as $V_{k}$-transverse knots, $(L_{1})_{i_{1},0}$ is isotopic to $(L_{1})_{0,i_{1}}^{i_{1}}$. In other words, up to $V_{k}$-transverse isotopy, $(L_{1})_{i_{1},0}$ can be obtained from $(L_{1})_{0,i_{1}}$ by adding $i_{1}$ pairs of kinks as in Figure  1. Similarly, up to $V_{k}$-transverse isotopy, $(L_{2})_{i_{2},0}$ can be obtained from and $(L_{2})_{0,i_{2}}$ by adding $i_{2}$ pairs of kinks as in Figure  1.

### v2

@@ 5. Examples
Let $V_{k}$ be a co-orienting vector fied of $\xi_{k}$. Let $K_{1}\sqcup K_{2}$ be a $V_{k}$-transverse link in $M$. Let $L=L_{1}\sqcup L_{2}$ be a Legendrian representative of the smooth isotopy class of $K_{1}\sqcup K_{2}$, obtained by taking a $C^{0}$-approximation, is in [ETN05]. Consider integers $i_{1},i_{2}\in\mathbb{Z}$. Recall from Section 2 that, as $V_{k}$-transverse knots, $(L_{1})_{i_{1},0}$ is isotopic to $(L_{1})_{0,i_{1}}^{i_{1}}$. In other words, up to $V_{k}$-transverse isotopy, $(L_{1})_{i_{1},0}$ can be obtained from $(L_{1})_{0,i_{1}}$ by adding $i_{1}$ pairs of kinks as in Figure  1. Similarly, up to $V_{k}$-transverse isotopy, $(L_{2})_{i_{2},0}$ can be obtained from and $(L_{2})_{0,i_{2}}$ by adding $i_{2}$ pairs of kinks as in Figure  1.